OpenStudy (zab505):

IL-1ß is a protein released by cells under conditions of infection. Excessive release of IL-1ß can be very damaging and can lead to severe disease. In a typical experiment, the typical concentration of IL-1ß around stressed cells is 800 ± 113 pg/ml. A researcher is testing drugs that can inhibit IL-1ß release. She tests two drugs, drug-1 and drug-2, for their effects on IL-1ß release. If the drugs have a z-score of -2.5, she will consider them promising for future study. In her experiment, stressed cells treated with drug-1 yielded 1104 pg/ml of IL-1ß, while drug-2 yielded 540 pg/ml of IL-1ß.

OpenStudy (zab505):

A. She will not study drug-2 or drug-1. B. She will continue studying drug-1 but not drug-2. C. She will continue studying both drug-1 and drug-2. D. She will continue studying drug-2 but not drug-1.

OpenStudy (zab505):

@ShailKumar

OpenStudy (zab505):

how is it going?

OpenStudy (anonymous):

I don't know how to calculate z-score. :(

OpenStudy (anonymous):

Practically it seems D

OpenStudy (zab505):

Thank you!

OpenStudy (tkhunny):

Calculating a z-score is a fundamental principle. You MUST know how to do this with accuracy and confidence. I will do one. You show me the other. You MUST get it and get it well. 800 ± 113 pg/ml. This is not quite a good enough bit of information. If this is intended to be the ENTIRE distribution, then 113 might be THREE standard deviations. I want to take it this way, but I believe this will not produce the right results. Let us PLAY LIKE 113 is the actual Standard Deviation of hte population. We have to assume something, or we cannot calculate a z-score. 1104 pg/ml (1104 - 800)/113 = 304/113 = +2.69 We're not interested in this drug AT ALL. 540 pg/ml Go!